OpenStudy (anonymous):

What is the sum of a 6–term geometric series if the first term is 6 and the last term is 966,306? :))

OpenStudy (anonymous):

Just use the formula for finding the sum of n-terms. It is :\[S _{n}=\frac{ n }{2 }(a _{1}+a _{n})\]

OpenStudy (anonymous):

You know n, you know a1, and you know an. Soo just plug away.

jimthompson5910 (jim_thompson5910):

calmat01, that formula only works for arithmetic series not geometric series

OpenStudy (anonymous):

Oh hell, I thought something was fishy about my reply. Arithmetic, not geometric.

OpenStudy (anonymous):

I think I deserve to be docked points. This is getting on my nerves.

jimthompson5910 (jim_thompson5910):

no it's ok

jimthompson5910 (jim_thompson5910):

The first thing you need to do is solve for r in the second equation below a(n) = a*r^(n - 1) 966306 = 6*r^(6 - 1) tell me what you get

OpenStudy (anonymous):

You are in good hands with jim, Hopeless, let him walk you through it. Disregard my posts.

jimthompson5910 (jim_thompson5910):

no the formulas you posted are still good, just for arithmetic series only

jimthompson5910 (jim_thompson5910):

anyways, what do you get when you solve 966306 = 6*r^(6 - 1) for r ?